a tank has a capacity of10 gal. when it is fill it contains 15% alcohol. how many gallons must be replaced by on 80% alcohol solution to give 10 gallons of 70% solution?
Answer by mananth(16946)

percent ----------------quantity gallons
80 ----------------x
15 ----------------10-x
70.00% ----------------10
>>>
80x+15(10-x)=70 *10
80x+150-15x= 700
80x-15x =700 - 150
65x=550
/65
x=8.46 gallons of 80%